OpenStudy (anonymous):

Evaluate: 32 + (6 − 2) • 4 − 6 over 3

OpenStudy (anonymous):

32 + (6 - 2) * 4 - 6/3 or is it .. (32 + (6 - 2) * 4 - 6) / 3

OpenStudy (anonymous):

need to know which one is the correct problem or you will not get the correct answer

OpenStudy (anonymous):

@kelliegirl33 i think its your second one, if it is then the answer is 14 i think.

OpenStudy (anonymous):

(32 + (6 - 2) * 4 - 6)/3 do parenthesis first (32 + 4 * 4 - 6)/3 do multiplication next (32 + 16 - 6)/3 add next (48 - 6) / 3 now subtract 42/3 now divide 14 I agree with you @thegeolance

OpenStudy (mathstudent55):

\( \dfrac{32 + (6 − 2) \cdot 4 − 6}{3} \) \( = \dfrac{32 + 4 \cdot 4 − 6}{3} \) \( = \dfrac{32 + 16 − 6}{3} \) \( = \dfrac{48 − 6}{3} \) \( = \dfrac{42}{3} \) \(= 14\)